# Rudin-Shapiro function along irreducible polynomials over finite fields

### v1

@@ Abstract.
Let $\mathbb{F}_{q}$ be the finite field of $q$ elements. We define the Rudin-Shapiro function $R$ on monic polynomials $f=t^{n}+f_{n-1}t^{n-1}+\dots+f_{0}\in\mathbb{F}_{q}[t]$ over $\mathbb{F}_{q}$ by
@@
We investigate the distribution of the Rudin-Shapiro function along the irreducible polynomials. We show, that the number of irreducible polynomials $f$ with $R(f)=\gamma$ for any $\gamma\in\mathbb{F}_{q}$ is asymptotically $q^{n-1}/n$ as $n\rightarrow\infty$.
@@ 1. Introduction
In recent years, many spectacular results have been obtained on important problems combining some arithmetic properties of the integers and some conditions on their digits in a given basis, see for example [2, 3, 4, 12, 13, 15, 18, 22, 26]. In particular, Mauduit and Rivat [13] showed that the Rudin-Shapiro sequence along prime numbers is well-distributed.
@@
Let $q$ be a prime power and $\mathbb{F}_{q}$ be the finite field of $q$ elements. For a monic polynomial $f\in\mathbb{F}_{q}[t]$ of degree $n\geq 2$, define the Rudin-Shapiro function $R(f)$ as
(1) $R(f)=\sum_{i=1}^{n-1}f_{i}f_{i-1}\quad\text{with }f=t^{n}+f_{n-1}t^{n-1}+\dots+f_{0}\in\mathbb{F}_{q}[t].$
@@
Our goal is to prove that the size of $R_{n}(\gamma)$ is asymptotically the same for all $\gamma$.
@@
#$\#{\mathcal{R}}_{n}(\gamma)=\frac{\#{\mathcal{P}}(n)}{q}+O\left(n^{3/2}q^{(1-\delta)n}\right),$
where the implied constant may depend only on $q$. One can choose $\delta=1/30$.
We remark, that for $r=2$ the result follows from the Weil bound, see comment before [17, Problem 8].
2. Auxiliary results
2.1. Notations
@@
Let us identify ${\mathcal{G}}(n)$ with the $n+1$ dimensional linear space $\mathbb{F}_{q}^{n+1}$ via
(3) $f=f_{n}t^{n}+\dots+f_{0}\in{\mathcal{G}}(n)\quad\longleftrightarrow\quad\mathbf{f}=(f_{0},\dots,f_{n})\in\mathbb{F}_{q}^{n+1}.$
Moreover, let us also identify $R$ with the $(n+1)\times(n+1)$ matrix $R=(R_{i,j})$ defined by
(4) $R_{i,j}=\left\{\begin{array}[]{cl}1&\text{if $i=j+1$}\\ 0&\text{otherwise}.\end{array}\right.$
Then we have $R(f)=\mathbf{f}^{T}R\mathbf{f}$, where $R(f)$ is defined as (1).
For given functions $F$ and $G$, the notations $F\ll G$, $G\gg F$ and $F=O(G)$ are all equivalent to the statement that the inequality $|F|\leq c|G|$ holds with some constant $c>0$. Throughout the paper, any implied constants in symbols $O$, $\ll$ and $\gg$ may depend only on $q$.
2.2. Quadratic forms of polynomials
Let $Q$ be a quadratic form on $\mathbb{F}_{q}^{n}$. Then, $Q$ can be written as $Q(x)=x^{T}Mx$ for some matrix $M\in\mathbb{F}_{q}^{n\times n}$. The rank of $Q$ is the rank of the matrix $M$. The first result we need is the following bound. For completeness, we also include the proof.
@@ 2.2. Quadratic forms of polynomials
Let $q$ be and odd prime power and let $Q$ be a quadratic and $L$ be a linear form on $\mathbb{F}_{q}^{n}$. Assume, that $Q$ has rank at least $r$. Then
$\left|\sum_{x\in\mathbb{F}_{q}^{n}}\psi(Q(x)+L(x))\right|\leq q^{n-r/2}$
for any non-trivial additive character $\psi$ of $\mathbb{F}_{q}$.
@@
Write $Q(x)=x^{T}Mx$ for some symmetric matrix $M\in\mathbb{F}_{q}^{n}$ of rank at least $r$ and let $B$ be the corresponding bilinear form
$B(x,y)=x^{T}My.$
Write also $B_{x}(y)=B(x,y)$. Let $K\leq\mathbb{F}_{q}^{n}$ of vectors $x$ such that $B_{x}=0$. It’s codimension is the rank of $M$ and thus of the quadratic form $Q$.
We have
$\left|\sum_{x\in\mathbb{F}_{q}^{n}}\psi(Q(x)+L(x))\right|^{2}$ $=\sum_{x,h\in\mathbb{F}_{q}^{n}}\psi(Q(x+h)+L(x+h))\overline{\psi(Q(x)+L(x))}$
$=\sum_{h\in\mathbb{F}_{q}^{n}}\psi(Q(h)+L(h))\sum_{x\in\mathbb{F}_{q}^{n}}\psi(2B_{h}(x)).$
If $h\not\in K$, then $B_{h}$ is a non-zero linear form and thus
$\sum_{x\in\mathbb{F}_{q}^{n}}\psi(2B_{h}(x))=\sum_{x\in\mathbb{F}_{q}^{n}}\psi(x)=0.$
For $h\in K$, we have
$\sum_{x\in\mathbb{F}_{q}^{n}}\psi(2B_{h}(x))=q^{n}.$
The assertion follows by noticing #$\#K=q^{n-r}$. ∎
Lemma 3.
Let $k<n/2$ and $a,b\in{\mathcal{M}}(k)$. For any fixed $b\in{\mathcal{M}}(k)$, let $r_{b}$ be the number of $a\in{\mathcal{M}}(k)$ such that the map $h\mapsto R(ah)-R(bh)$ with $h\in{\mathcal{M}}(n-k)$ has rank at most $n-2k$. Then $r_{b}=O(q^{(k+1)/2})$, where the implied constant is absolute.
Proof.
For a given polynomial $f\in{\mathcal{G}}(k)$, write $[f]\in\mathbb{F}_{q}^{(n+1)\times(n-k+1)}$ be the matrix
$[f]_{i,j}=\left\{\begin{array}[]{cl}f_{i-j}&\text{if $i\leq j$}\\ 0&\text{otherwise}.\end{array}\right.$
Observe, that if $g\in{\mathcal{G}}(n-k)$, then writing $\mathbf{g}=(g_{0},\dots,g_{n-k})^{T}$, $[f]\mathbf{g}$ corresponds to the product polynomial $f\cdot g$ via (3). It yields $R(fg)=\mathbf{g}^{T}[f]^{T}R[f]\mathbf{g}$, where the matrix $R$ is defined as (4). Then the map $h\mapsto R(ah)-R(bh)$ corresponds to $M_{a,b}=[a]^{T}R[a]-[b]^{T}R[b]$.
First write $A=[a]^{T}R[a]\in\mathbb{F}_{q}^{(n-k+1)\times(n-k+1)}$. Observe, that the entries of $A$ at position $(i,j)$ with $i\leq j\leq n-k$ are
$A_{i,j}=\sum_{\ell=0}^{k-j+i-1}a_{\ell}a_{\ell+j-i+1}.$
Specially, $A_{i,j}=0$ for $j>i+k$.
Define $S=(s_{i,j})\in\mathbb{F}_{q}^{(k+1)\times(k+1)}$ as
$s_{i,j}=\left\{\begin{array}[]{cl}1&\text{if $i=j+1$}\\ 0&\text{otherwise}.\end{array}\right.$
Then
$A_{i,j}=\mathbf{a}^{T}S^{j-i+1}\mathbf{a},\quad\text{where }\mathbf{a}=(a_{0},\dots,a_{k})^{T}\in\mathbb{F}_{q}^{k+1}.$
Similarly, writing $b=[b]^{T}R[b]\in\mathbb{F}_{q}^{(n-k+1)\times(n-k+1)}$, then
$B_{i,j}=\mathbf{b}^{T}S^{j-i+1}\mathbf{b},\quad\text{where }\mathbf{b}=(a_{0},\dots,b_{k})^{T}\in\mathbb{F}_{q}^{k+1}.$
It yields,
$(M_{a,b})_{i,i+d}=\left\{\begin{array}[]{cc}\mathbf{a}^{T}S^{d+1}\mathbf{a}-\mathbf{b}^{T}S^{d+1}\mathbf{b}&\text{if $0\leq d<k$}\\ 0&\text{if $d\geq k$}.\end{array}\right.$
Assume, that $\mathbf{a}$ is so that
(5) $\mathbf{a}^{T}S^{d+1}\mathbf{a}-\mathbf{b}^{T}S^{d+1}\mathbf{b}\neq 0$
for some $0\leq d<k$, and let $d^{*}$ be the maximal index such that (5) holds with $d^{*}$ instead if $d$. Then the submatrix of $(M_{a,b})$ formed by $(M_{a,b})_{i,j}$ with $i=1,\dots,n-2k+1$ and $j=d^{*},\dots,d^{*}+n-2k+1$ is a lower-triangular matrix with non-zero elements in the diagonal, thus it has rank $n-2k+1$.
Thus, to prove the result, it is enough to estimate the number $N$ of $a\in{\mathcal{M}}(k)$ such that (5) does not hold for any $0\leq d<k$, that is
$\mathbf{a}^{T}S^{d+1}\mathbf{a}=\mathbf{b}^{T}S^{d+1}\mathbf{b}\quad 0\leq d<k.$
Let $\psi$ be a non-trivial additive character of $\mathbb{F}_{q}$. Then
$N$ $=\frac{1}{q^{k}}\sum_{a\in{\mathcal{M}}(k)}\sum_{t_{1},\dots,t_{k}}\psi\left(\sum_{i=1}^{k}t_{i}\left(\mathbf{a}^{T}S^{i}\mathbf{a}-\mathbf{b}^{T}S^{i}\mathbf{b}\right)\right)$
$=\frac{1}{q^{k}}\sum_{t_{1},\dots,t_{k}}\psi\left(-\mathbf{b}^{T}\left(\sum_{i=1}^{t}t_{i}S^{i}\right)\mathbf{b}\right)\sum_{a\in{\mathcal{M}}(k)}\psi\left(\mathbf{a}^{T}\left(\sum_{i=1}^{k}t_{i}S^{i}\right)\mathbf{a}\right).$
The entries of
(6) $t_{1}S+t_{2}S^{2}+\dots+t_{k}S^{k}\in\mathbb{F}_{q}^{(k+1)\times(k+1)}$
at position $(i,i+u)$ is $t_{u}$ for $1\leq u\leq k$ and 00 for $u>k$. Thus the rank of (6) is $k+1-\min\{1\leq u\leq k:t_{u}\neq 0\}$.
As the rank of (6) is reduced at most one if we consider it on the affine space of monic polynomials $a$, we get by Lemma 2,
$N$ $\leq 1+\frac{1}{q^{k}}\sum_{u=1}^{k}\sum_{\begin{subarray}{c}t_{v}\in\mathbb{F}_{q}\\ u<v\leq k\end{subarray}}\sum_{t_{u}\in\mathbb{F}_{q}^{*}}\left|\sum_{a\in{\mathcal{M}}(k)}\psi\left(\mathbf{a}^{T}\left(\sum_{i=u}^{k}t_{i}S^{i}\right)\mathbf{a}\right)\right|$
$\ll\frac{1}{q^{k}}\sum_{u=1}^{k}q^{k-u+1}q^{k-(k-u)/2}\ll q^{(k+1)/2}.$
Lemma 4.
Let $a\in{\mathcal{M}}(k)$. Then for $h\in{\mathcal{M}}(n-k)$, the bilinear map $h\mapsto R(ah)$ has rank at least $n-k-1$.
Proof.
Write $a=t^{u}b$ with $t\nmid b$ and $b=t^{k-u}+b_{k-u-1}t^{k-u-1}+\dots+b_{0}$. Moreover, define $B=(B_{i,j})\in\mathbb{F}_{q}^{(n-u)\times(n-k)}$ as
$B_{i,j}=\left\{\begin{array}[]{cl}b_{i-j}&\text{if $i\geq j$}\\ 0&\text{otherwise},\end{array}\right.$
and $S=(S_{i,j})\in\mathbb{F}_{q}^{(n-u)\times(n-u)}$ as
$S_{i,j}=\left\{\begin{array}[]{cl}1&\text{if $i=j+1$}\\ 0&\text{otherwise}.\end{array}\right.$
Then, the bilinear map $R(ah)$ corresponds to the matrix $B^{T}SB$.
Write
$B=\left(\begin{matrix}B^{\prime}\\ P\end{matrix}\right)\quad\text{and}\quad S=\left(\begin{matrix}S_{1}&\mathbf{0}\\ S_{2}&S_{3}\end{matrix}\right)$
with $B^{\prime}\in\mathbb{F}_{q}^{(n-k)\times(n-k)}$, $P\in\mathbb{F}_{q}^{(k-u)\times(n-k)}$, $S_{1}\in\mathbb{F}_{q}^{(n-k)\times(n-k)}$, $S_{2}\in\mathbb{F}_{q}^{(k-u)\times(n-k)}$ and $S_{3}\in\mathbb{F}_{q}^{(k-u)\times(k-u)}$.
As $t\nmid b$, $b_{0}\neq 0$, so $B^{\prime}$ is a lower triangular matrix with non-zero element in the diagonal, thus has full rank $n-k$. Also, $S_{1}$ has rank $n-k-1$. Thus
$B^{T}SB=\left(\begin{matrix}(B^{\prime})^{T}S_{1}B^{\prime}&\mathbf{0}\\ P^{T}S_{2}B^{\prime}&P^{T}S_{3}P\end{matrix}\right)$
has the submatrix $(B^{\prime})^{T}S_{1}B^{\prime}$ of rank $n-k-1$. ∎
2.3. Divisor bounds
Let $\tau(f)$ denote the number of monic divisors of $f\in\mathbb{F}_{q}[t]$. We need the following bound on the second moment of $\tau$, see [1, Lemma 8].

### v2

@@ Abstract.
Let $q$ be an odd prime power and $\mathbb{F}_{q}$ be the finite field of $q$ elements. We define the Rudin-Shapiro function $R$ on monic polynomials $f=t^{n}+f_{n-1}t^{n-1}+\dots+f_{0}\in\mathbb{F}_{q}[t]$ over $\mathbb{F}_{q}$ by
@@
We investigate the distribution of the Rudin-Shapiro function along irreducible polynomials. We show that the number of irreducible polynomials $f$ with $R(f)=\gamma$ for any $\gamma\in\mathbb{F}_{q}$ is asymptotically $q^{n-1}/n$ as $n\rightarrow\infty$.
@@ 1. Introduction
In recent years, many spectacular results have been obtained on important problems combining some arithmetic properties of integers and some conditions on their digits in a given basis, see for example [2, 3, 4, 12, 13, 15, 18, 22, 26]. In particular, Mauduit and Rivat [13] showed that the Rudin-Shapiro sequence along prime numbers is well-distributed.
@@
Let $q$ be an odd prime power and $\mathbb{F}_{q}$ be the finite field of $q$ elements. For a monic polynomial $f\in\mathbb{F}_{q}[t]$ of degree $n\geq 2$, define the Rudin-Shapiro function $R(f)$ as
$R(f)=\sum_{i=1}^{n-1}f_{i}f_{i-1}\quad\text{with }f=t^{n}+f_{n-1}t^{n-1}+\dots+f_{0}\in\mathbb{F}_{q}[t].$
@@
Our goal is to prove that the size of ${\mathcal{R}}_{n}(\gamma)$ is asymptotically the same for all $\gamma$.
@@
${\mathcal{R}}_{n}(\gamma)=\frac{\#{\mathcal{P}}(n)}{q}+O\left(n^{3/2}q^{(1-\delta)n+1}\right),$
where the implied constant is absolute. One can choose $\delta=1/28$.
We remark, that for $n=2$ the result follows from the Weil bound, see comment before [17, Problem 8].
By Theorem 1, we have that
${\mathcal{R}}_{n}(\gamma)\sim\frac{\#{\mathcal{P}}(n)}{q}$
if either $n\rightarrow\infty$ or $q\rightarrow\infty$ provided $n>2/\delta$.
2. Outline of the proof
@@
For $\ell\geq 0$, write
$S^{(\ell)}(f)=\sum_{i=\ell}^{n}f_{i}f_{i-\ell}\quad\text{with }f=f_{n}t^{n}+f_{n-1}t^{n-1}+\dots+f_{0}\in\mathbb{F}_{q}[t].$
For $\ell=1$, we simply write
$S(f)=S^{(1)}(f).$
Clearly, for monic polynomials $f$ we have
(2) $R(f)=S(f)-f_{n-1}.$
In order to prove Theorem 1, we introduce Vaughan’s identity for polynomials in Section 4.4. Then the main tool is to investigate single and double exponential sums in terms of $S(f)$ in Section 4.1. The bounds of the exponential sum estimates are given in terms of the rank of quadratic forms which is studied in Sections 4.2 and 4.3. Then we conclude the proof of Theorem 1 in Section 5.
For given functions $F$ and $G$, the notations $F\ll G$, $G\gg F$ and $F=O(G)$ are all equivalent to the statement that the inequality $|F|\leq c|G|$ holds with some constant $c>0$.
3. preliminaries
Let $\tau(f)$ denote the number of monic divisors of $f\in\mathbb{F}_{q}[t]$. First, we state the following bound on $\tau(f)$.
@@
For any $\varepsilon>0$ we have
$\tau(f)\ll\min\left\{|f|^{(2+\varepsilon)/\log\deg f},2^{\deg f}\right\},$
where the implied constant can depend on $\varepsilon$.
@@
The first bound follows from [16, Lemma 16], while the second bound follows from the fact that a polynomial $f$ has at most $\deg f$ irreducible divisor. ∎
We also have the following bound on the second moment of $\tau$, see [1, Lemma 8].
@@
4. Auxiliary results
4.1. Exponential sums of quadratic forms of polynomials
Let $Q$ be a quadratic form on $\mathbb{F}_{q}^{n}$. Then, $Q$ can be written as $Q({\mathbf{x}})={\mathbf{x}}^{T}M{\mathbf{x}}$ for some symmetric matrix $M\in\mathbb{F}_{q}^{n\times n}$. The rank of $Q$ is the rank of the matrix $M$. The first result we need is the following bound. For completeness, we also include the proof.
Lemma 4.
Let $q$ be an odd prime power and let $Q$ be a quadratic form, and let $L$ be a linear form on $\mathbb{F}_{q}^{n}$. Assume that $Q$ has rank at least $r$. Then
$\left|\sum_{{\mathbf{x}}\in\mathbb{F}_{q}^{n}}\psi(Q({\mathbf{x}})+L({\mathbf{x}}))\right|\leq q^{n-r/2}$
for any non-trivial additive character $\psi$ of $\mathbb{F}_{q}$.
Proof.
Write $Q({\mathbf{x}})={\mathbf{x}}^{T}M{\mathbf{x}}$ for some symmetric matrix $M\in\mathbb{F}_{q}^{n\times n}$ of rank at least $r$ and let $B$ be the corresponding bilinear form
$B({\mathbf{x}},{\mathbf{y}})={\mathbf{x}}^{T}M{\mathbf{y}}.$
Write also $B_{\mathbf{x}}({\mathbf{y}})=B({\mathbf{x}},{\mathbf{y}})$. Let $K\leq\mathbb{F}_{q}^{n}$ the subspace of vectors $\mathbf{x}$} such that $B_{\mathbf{x}}=0$. It’s codimension is the rank of $M$ and thus of the quadratic form $Q$.
We have
$\left|\sum_{{\mathbf{x}}\in\mathbb{F}_{q}^{n}}\psi(Q({\mathbf{x}})+L({\mathbf{x}}))\right|^{2}$
$=\sum_{{\mathbf{x}},{\mathbf{h}}\in\mathbb{F}_{q}^{n}}\psi(Q({\mathbf{x}}+{\mathbf{h}})+L({\mathbf{x}}+{\mathbf{h}}))\overline{\psi(Q({\mathbf{x}})+L({\mathbf{x}}))}$
$=\sum_{{\mathbf{h}}\in\mathbb{F}_{q}^{n}}\psi(Q({\mathbf{h}})+L({\mathbf{h}}))\sum_{{\mathbf{x}}\in\mathbb{F}_{q}^{n}}\psi(2B_{\mathbf{h}}({\mathbf{x}})).$
If ${\mathbf{h}}\not\in K$, then $B_{\mathbf{h}}$ is a non-zero linear form and thus
$\sum_{{\mathbf{x}}\in\mathbb{F}_{q}^{n}}\psi(2B_{h}({\mathbf{x}}))=\sum_{{\mathbf{x}}\in\mathbb{F}_{q}^{n}}\psi({\mathbf{x}})=0.$
For ${\mathbf{h}}\in K$, we have
$\sum_{{\mathbf{x}}\in\mathbb{F}_{q}^{n}}\psi(2B_{\mathbf{h}}({\mathbf{x}}))=q^{n}.$
The assertion follows by noticing $K=q^{n-r}$. ∎
4.2. Reversed polynomials
For a polynomial $a\in\mathbb{F}_{q}[t]$ of form $a=a_{n}t^{n}+\dots+a_{1}t+a_{0}$ let $a^{*}$ be the polynomial with reversed coefficients, that is,
$a^{*}(t)=t^{n}a(t^{-1})=a_{0}t^{n}+\dots+a_{n-1}t+a_{n}\in\mathbb{F}_{q}[t].$
First we show, that the quantities $S^{(\ell)}(a)$ can be determined by using the reversed polynomial $a^{*}$ of $a$.
Lemma 5.
For any polynomial $a\in{\mathcal{G}}(n)$, we have
$a^{*}\cdot a=\sum_{\ell=0}^{n}S^{(\ell)}(a)\left(t^{n+\ell}+t^{n-\ell}\right).$
Proof.
The coefficient of the term $t^{n-\ell}$ in $a^{*}\cdot a$ is
$\sum_{i+j=n-\ell}a_{n-i}a_{j}=\sum_{j=0}^{n-\ell}a_{j+\ell}a_{j}=S^{(\ell)}(a).$
Moreover, the coefficient of $t^{n-\ell}$ is the same as the coefficient of $t^{n+\ell}$. ∎
Next, we investigate the following type of quadratic equations in terms of the polynomials $a$ and its reversed $a^{*}$.
Lemma 6.
Let $\varepsilon>0$. For given integer $n$, and polynomial $f\in{\mathcal{G}}(2n)$, let $N(f)$ be the number of polynomials $a\in{\mathcal{G}}(n)$ with
(3) $a^{*}\cdot a=f.$
Then
$N(f)\ll\min\left\{2^{n},q^{(2+\varepsilon)\frac{n}{\log_{q}n}}\right\},$
where the implied constant may depend on $\varepsilon$.
Proof.
Clearly, for any polynomials $a,b\in\mathbb{F}_{q}[t]$, we have
(4) $(a\cdot b)^{*}=a^{*}\cdot b^{*}.$
Now assume that $f$ has the form
$f=b^{*}\cdot b\quad\text{for some }b\in{\mathcal{G}}(n),$
since otherwise the bound trivially holds.
If $a$ is a polynomial satisfying (3), put
$d=\gcd(a,b).$
Then by (4)
$f=d\cdot\frac{b}{d}\cdot d^{*}\cdot\left(\frac{b}{d}\right)^{*},$
and as
$\gcd(a/d,b/d)=1\quad\text{and}\quad\gcd((a/d)^{*},(b/d)^{*})=1$
we must have
$a=d\cdot\left(\frac{b}{d}\right)^{*}.$
That is, the divisor $d$ uniquely determines $a$. Whence, the number of polynomials $a$ with (3) is at most $\tau(b)$. Then the bound follows from Lemma 2. ∎
4.3. Rank of quadratic forms
For a fixed $a\in{\mathcal{M}}(k)$ the map $h\mapsto R(ah)$ with $h\in{\mathcal{G}}(n-k)$ is a quadratic form in the coefficients of $h$. First we investigate its rank.
Lemma 7.
Let $a\in{\mathcal{M}}(k)$ and for $h\in{\mathcal{G}}(n-k)$ consider the map $Q_{a}:h\mapsto S(ah)$, which is a quadratic form in the coefficients of $h$. If $k<n/2$, then it has rank at least $\mathop{\mathrm{rank}}Q_{a}\geq n-k-1$.
Proof.
Let $B$ be the symmetric bilinear form so that $S(h)=B(h,h)$. Specially,
(5) $B(t^{i},t^{j})=\left\{\begin{array}[]{cl}1/2&\text{if }|i-j|=1,\\ 0&\text{otherwise},\end{array}\right.\quad 0\leq i,j\leq n.$
Put
$K=\{g\in{\mathcal{G}}(n-k):h\mapsto B(ag,ah)\text{ is the zero map}\}$
which is a linear subspace of ${\mathcal{G}}(n-k)$. Then
(6) $\mathop{\mathrm{rank}}Q_{a}=n-k+1-\dim K.$
Write
$r^{(i)}\equiv t^{i}\mod a,\quad\deg r^{(i)}<\deg a$
and
$b^{(i)}=\left\{\begin{array}[]{cl}t^{i}&\text{if }i<k,\\ t^{i}-r^{(i)}&\text{otherwise},\end{array}\right.\quad i=0,1,\dots,n.$
Clearly, for each polynomial $f\in{\mathcal{G}}(n)$ of form
$f=\sum_{i=0}^{n}f_{i}t^{i}$
one has
$f=\sum_{i=0}^{n}c_{i}b^{(i)}\quad\text{ for some }c_{0},c_{1},\dots,$
moreover
(7) $a\mid f\quad\Leftrightarrow\quad c_{0}=\dots=c_{k-1}=0,\text{ and }c_{k}=f_{k},\dots,c_{n}=f_{n}.$
Indeed, if $a\mid f$, then
$f=f-(f\bmod a)=\sum_{i=0}^{n}f_{i}(t^{i}-r^{(i)})=\sum_{i=k}^{n}f_{i}b^{(i)},$
as $t^{i}=r^{(i)}$ for $0\leq i<k$. Hence, $b^{(k)},\dots,b^{(n)}$ is a basis of the vector space of polynomials of form $ah$, $h\in{\mathcal{G}}(n-k+1)$.
Now assume, that $g\in K$ and write
$ag=\sum_{i=0}^{n}d_{i}t^{i}.$
By (7), we can also write
(8) $ag=\sum_{i=k}^{n}d_{i}b^{(i)}.$
As $g\in K$, we have by (5) that for $\ell=k,\dots,n$
$0$ $=B(ag,b^{(\ell)})=B\left(\sum_{i=0}^{n}d_{i}t^{i},t^{\ell}-r^{(\ell)}\right)$
(9) $=B\left(\sum_{i=0}^{n}d_{i}t^{i},t^{\ell}\right)-B\left(\sum_{i=0}^{n}d_{i}t^{i},r^{(\ell)}\right)$
$=\frac{d_{\ell+1}+d_{\ell-1}}{2}-B\left(\sum_{i=0}^{n}d_{i}t^{i},r^{(\ell)}\right).$
As $\deg r^{(\ell)}<k$, we have $B(t^{i},r^{(\ell)})=0$ for $i>k$. The (8) and (4.3) yield that
$d_{\ell+1}+d_{\ell-1}=2B\left(\sum_{i=0}^{n}d_{i}t^{i},r^{(\ell)}\right)=2d_{k}B\left(r^{(k)},r^{(\ell)}\right),\quad\text{for }k\leq\ell<n.$
That is, $d_{k}$ and $d_{k+1}$ determine all the values of $d_{k+2},\dots,d_{n}$. Also, as
$\sum_{i=0}^{k-1}d_{i}t^{i}=\sum_{i=k}^{n}d_{i}(t^{i}\bmod a),$
the values $d_{0},\dots,d_{k-1}$ are also uniquely determined. Thus $\dim K=2$. Then the result follows from (6). ∎
In the next lemma we investigate the rank of the quadratic form $h\mapsto S(ah)-S(bh)$ with $h\in{\mathcal{G}}(n-k)$.
Lemma 8.
Let $a,b\in{\mathcal{M}}(k)$ be polynomials such that
(10) $a^{*}\cdot a\neq b^{*}\cdot b.$
Then, the quadratic form $h\mapsto S(ah)-S(bh)$ with $h\in{\mathcal{G}}(n-k)$ has rank at least n2k1n-2k-1.
Proof.
Recall, that the symmetric bilinear form $B$ defined by (5) satisfies $S(h)=B(h,h)$.
First consider the symmetric bilinear form $B_{a}$ defined by
$B_{a}(h,g)=B(ah,ag),$
with $h,g\in{\mathcal{G}}(n-k)$. Then
$B_{a}(t^{i},t^{j})=B(a\cdot t^{i},a\cdot t^{j})=\frac{1}{2}\left(S^{(|i-j-1|)}(a)+S^{(|j-i-1|)}(a)\right).$
Similarly, if $B_{b}$ is the symmetric bilinear form defined by $B_{b}(h,g)=B(bh,bg)$, $h,g\in{\mathcal{G}}(n-k)$, we also have
$B_{b}(t^{i},t^{j})=B(b\cdot t^{i},b\cdot t^{j})=\frac{1}{2}\left(S^{(|i-j-1|)}(b)+S^{(|j-i-1|)}(b)\right).$
Whence, the symmetric bilinear form $B_{a,b}$ corresponding to S(ah)S(bh)S(ah)-S(bh) satisfies
(11) $B_{a,b}(t^{i},t^{j})=\frac{1}{2}\left(S^{(|i-j-1|)}(a)+S^{(|j-i-1|)}(a)-S^{(|i-j-1|)}(b)-S^{(|j-i-1|)}(b)\right).$
Let $M\in\mathbb{F}_{q}^{(n-k+1)\times(n-k+1)}$ be the symmetric matrix corresponding to $B_{a,b}$ with respect to the basis $1,t,t^{2},\dots,t^{n-k}$. We show, that it has rank at least n2k1n-2k-1. To do so let $\ell$ be the smallest integer such that $t^{\ell}\nmid a^{*}\cdot a-b^{*}\cdot b$. By (10), $\ell$ is well-defined. Also, as $a^{*}\cdot a-b^{*}\cdot b$ is symmetric, $\ell\leq k$. Then, by the definition of $\ell$ and by Lemma 5 we have
(12) $S^{(\nu)}(a)=S^{(\nu)}(b)\quad\text{for }k-\ell<\nu\leq k\quad\text{and}\quad S^{(k-\ell)}(a)\neq S^{(k-\ell)}(b).$
By definition of $S^{(\nu)}$, we also have, that
(13) $S^{(\nu)}(a)=S^{(\nu)}(b)=0\quad\text{for }\nu>k.$
Then the submatrix
$(M_{i,j})\quad\text{with }k-\ell+2\leq i\leq n-k+1,\quad 1\leq j\leq n-2k+\ell$
of $M$ is an upper triangular matrix with nonzero elements in the diagonal by (12) and (11). Indeed, for $0\leq\nu,\mu<n-2k+\ell$ with $\nu>\mu$, we have
$M_{k-\ell+2+\nu,1+\mu}=\frac{1}{2}\Big{(}$ $S^{(|k-\ell+\nu-\mu|)}(a)+S^{(|\mu-\nu-k+\ell-2|)}(a)$
$\quad-S^{(|k-\ell+\nu-\mu|)}(b)-S^{(|\mu-\nu-k+\ell-2|)}(b)\Big{)}=0.$
by (13) as $|k-\ell+\nu-\mu|>k$-$\ell$ and $|\mu-\nu-k+\ell-2|>k-\ell+2$. Moreover, the values in the diagonal are
$M_{k-\ell+2+\nu,1+\nu}$
$=\frac{1}{2}\left(S^{(|k-\ell|)}(a)+S^{(|-k+\ell-2|)}(a)-S^{(|k-\ell|)}(b)-S^{(|-k+\ell-2|)}(b)\right)$
$=\frac{1}{2}\left(S^{(k-\ell)}(a)-S^{(k-\ell)}(b)\right)\neq 0$
by (12) and (13) as $|-k+\ell-2|>k$-$\ell$.
Thus it has rank $n-2k+\ell-1$ and so
$\mathop{\mathrm{rank}}M\geq n-2k+\ell-1\geq n-2k-1.$
which proves the result. ∎